Sue rowed her boat across lake bend and back in 3hrs. If her rate returning was 2 mph less than the rate going, and if the distance each way was 7 miles, find her rate going.
Answer by stanbon(75887)

Sue rowed her boat across lake bend and back in 3hrs. If her rate returning was 2 mph less than the rate going, and if the distance each way was 7 miles, find her rate going.
------------
Going DATA:
rate = x mph ; distance = 7 miles ; time = d/r = 7/x hrs
-----------------------
Coming DATA:
rate = "x-2" mph ; distance = 7 miles ;time = d/r = 7/(x-2)
-------------------------------
Equation:
time + time = 3 hrs
7/x + 7/(x-2) = 3
Multiply thru by x(x-2):
7(x-2) + 7x = 3x(x-2)
7x-14 + 7x = 3x^2-6x
3x^2 - 20x + 14 = 0
x = 5.872 mph (this is her going rate)
========================================
Cheers,
Stan H.